OpenStudy (anonymous):

solve: cos^(2x)-sin(2x)=0

OpenStudy (anonymous):

please clarify cos^(2x) and please give me a medal for the last problem :P

OpenStudy (anonymous):

my b, cos^2(2x)-sin^2(2x)=0

OpenStudy (anonymous):

\[2x=\theta \] \[\cos^2(\theta)-\sin^2(\theta)=0 \iff \cos^2(\theta)=\sin^2(\theta)\] \[\rightarrow \cos(\theta)=\pm \sin(\theta)\] \[\theta=\frac{k \pi}{4},k \in \mathbb{Z} \iff 2x=\frac{k \pi}{4}, k \in \mathbb{Z} \iff x=\frac{k \pi}{8}, k \in \mathbb{Z}\]

OpenStudy (anonymous):

x=2npi-pi/4 n is an integer

OpenStudy (anonymous):

\[\forall n \in \mathbb{Z}, 2 \pi n=0[\mod2\pi]\] So there's no need for the 2npi